OpenStudy (anonymous):

Solve the following system. y = x + 3 2x + y = 9

OpenStudy (anonymous):

2x+x+3=9 put value of y here. 3x=6 x=2 y=2+3=5

OpenStudy (hoblos):

replace the first equation in the second 2x + x+3 = 9 3x = 6 x=2 y=x+3=2+3=5